OpenStudy (anonymous):

what percent of 70 is 21?

OpenStudy (savvy):

30%

OpenStudy (anonymous):

x = 21 * 100 / 70 = 210/7 = 30%

OpenStudy (anonymous):

21/70, put it in a calculator...if you don't have one. SImplify the fraction 21/70 = 3/10 (both have 7 in)